OpenStudy (anonymous):

if the poly.x^4-6x*+16x^2-25x+10 is divided by another poly.x^2-2x+k,the remainder comes out to be x+a,find k and a

sam (.sam.):

what's that star * in .x^4-6x*+16x^2-25x+10

OpenStudy (anonymous):

cube

sam (.sam.):

x^4-6x*+16x^2-25x+10 Doing synthetic division you'll get the remainder (-9+2k)x+10-8k+k^2

sam (.sam.):

given remainder = x + a so, x+a = (- 9 + 2k)x + 10 – 8k + k Comparing coefficients of x on both sides 1 = - 9 + 2k or 2k = 10 or k = 5 Comparing constants on both sides 10 – 8k + k^2=a 10-8(5)+5^2=a a=-5

OpenStudy (lgbasallote):

hypothetical question @parvathysubhash the denotation of ^ as square and * as cube were given to you by your teacher?